OpenStudy (anonymous):

im making cookies but i only get 30 cookies and im making cookies for 50 people how do i make a proportion for all the ingredients so that i can make enough cookies?

OpenStudy (anonymous):

im using 225 grams of butter but that's enough for 30 cookies, what equation do i do to make it enough for 50?

OpenStudy (mathmale):

Personally I'd just double the recipe! 50 is 5/3 of 30, so you could multiply every measurement by 5/3 to get enough dough for 50 cookies.

OpenStudy (anonymous):

\[\frac{ x }{ 225 }*\frac{ 5 }{3 }\]

OpenStudy (anonymous):

would this work?

OpenStudy (mathmale):

I have no way of knowing where your fraction x/225 came from. Please elaborate.

OpenStudy (anonymous):

well... i'm trying to find how many grams of butter i would need to make 50 cookies and I already know that 225g is enough for 30 cookies.

OpenStudy (anonymous):

therefore \[\frac{ x }{ 225 }=\frac{ 5 }{3}\]